### v3

# Student Member, IEEE

Abstract
Continuously optimizing sensor placement is essential for precise target localization in various military and civilian applications. While information theory has shown promise in optimizing sensor placement, many studies oversimplify sensor measurement models or neglect dynamic constraints of mobile sensors. To address these challenges, we employ a range measurement model that incorporates radar parameters and radar-target distance, coupled with Model Predictive Path Integral (MPPI) control to manage complex environmental obstacles and dynamic constraints. We compare the proposed approach against stationary radars or simplified range measurement models based on the Root Mean Squared Error (RMSE) of the Cubature Kalman Filter (CKF) estimator for the targets’ state. Additionally, we visualize the evolving geometry of radars and targets over time, highlighting areas of highest measurement information gain, demonstrating the strengths of the approach. The proposed strategy outperforms stationary radars and simplified range measurement models in target localization, achieving a 38-74% reduction in mean RMSE and a 33-79% reduction in the upper tail of the 90% Highest Density Interval (HDI) over 500 Monte Carlo (MC) trials across all time steps. Code available at https://github.com/mlpotter/OptimalTrackingAndControl.
{IEEEkeywords}
Model Predictive Path Integral, Model Predictive Control, Radar, Cubature Kalman Filter
1 INTRODUCTION
Deploying a fleet of autonomous mobile sensors to estimate the positions and velocities of targets over time is an important research problem, with applications in transportation, navigation, defense, surveillance, and emergency rescue [1, 2, 3, 4], exemplified in Figure 1. This problem involves multiple challenges, including predicting expected sensor information gain without exact target or measurement data, adapting mobile sensors planned trajectories to dynamic environmental and target changes, and selecting suitable sensor models and modalities. We will discuss these challenges in the following paragraphs.
The multi-target tracking and sensor placement literature has explored various sensing modalities, including cameras [5], ultrasonic sensors [6, 7], Light Detection and Ranging (LIDAR) [8], acoustic range sensors [9], and radar-based techniques dating back to World War II. This paper focuses on tracking with radar in the Ultra High Frequency (UHF) (particularly “cognitive radars” [10]), due to the long-range detection capability, ability to penetrate atmospheric obstacles, and versatility in capturing detailed target features [11].
(a) Search and rescue in forest fire
(b) Enemy drone localization and tracking
Figure 1: Subfigure (a) shows the civilian application of a search and rescue mission where mobile radars navigate through a dense forest (on fire) to quickly localize and track lost victims. Subfigure (b) shows the military application of enemy drone localization and tracking in a city environment during a cloudy day.
A key challenge in radar-based target tracking is in incorporating realistic sensor models into the information metrics used for controlling the sensors. In particular, most existing control strategies for this problem use the Posterior Fisher Information Matrix (PFIM) [12] (inverse of the Posterior Cramer Rao Lower Bound (PCRLB)) to quantify the relationship between radar-target geometry and target localization performance [13, 14, 15]. However, studies optimizing sensor placement often oversimplify by using a constant covariance matrix for range measurements, neglecting signal path attenuation effects [4]. The approach in [4] focuses on optimizing the angle subtended between the radars and targets but does not encourage the mobile radars to move closer to the target. In [16], the authors provide the error covariance matrix of a Best Linear Unbiased Estimator (BLUE) estimator for the range measurement using a first-order Taylor Series approximation of a distance dependent range (DDR) measurement model, which yields the Fisher Information Matrix (FIM). Similarly, we calculate the FIM for the DDR measurement model from [16], where the covariance matrix of the range measurement is directly proportional to the received power at the radar after signal path attenuation, but we do not apply a first-order Taylor Series approximation to the DDR measurement model. This strategy encourages mobile radars to move closer to the targets, thereby increasing the Signal Noise Ratio (SNR) and improving radar range measurements.
Another major challenge in scenarios with moving targets and movable radars is that the optimal radar-target geometry may change substantially over time. Consequently, many studies use Model Predictive Controller (MPC) [17] for choosing the control inputs of mobile sensors over a finite horizon in response to uncertain target state estimates [18, 19, 20, 21, 22]. However, natural choices for the MPC objective in this problem, such as minimizing/maximizing the D-optimality of PCRLB/PFIM, lead to challenging nonlinear optimization problems that are computationally prohibitive to solve exactly. To address this issue, we instead utilize Model Predictive Path Integral (MPPI) to approximate the optimal control input distribution for the mobile radars. MPPI enables the handling of discontinuous, non-smooth objectives through highly parallelized computation for real-time control [23, 24]. However, as the number of agents, planning horizon, and the actuation space increases, parameter estimation, particularly for the covariance, requires an exponentially growing number of samples. To mitigate this issue and reduce the number of proposal distribution updates and trajectory samples needed in the high-dimensional control space, we integrate advanced Adaptive Importance Sampling (AIS) techniques, such as cross-entropy weighting with covariance shrinkage [25], to prevent overfitting of the sample covariance matrix.
This paper introduces a improved pipeline for continuously optimizing radar placement, aiming to overcome the previously mentioned challenges. We focus on range-only radars, deriving objective functions using FIMs that incorporate signal path attenuation in the range measurement model. These functions are optimized with respect to radar control inputs using MPPI control. Crucially, MPPI control, which is rapidly gaining popularity in control research, offers enhanced flexibility for real-time optimization of nonlinear, discontinuous, and non-differentiable objectives. These advantages enable MPPI to handle future extensions of the proposed objective, such as incorporating more complex signal measurement models or adding more constraints through occupancy grid maps.
The paper is structured as follows: Section 2 describes the current literature; Section 3 describes the simulation assumptions and notation; Section 4 describes the radar signal model and DDR measurement model; Section 5 formulates the targets’ state-space model and the radars’ kinematic model; Section 6 derives the FIM which accounts for the radar signal attenuation; Section 7 formulates our objective function based on the derived FIMs and outlines our MPPI controller, and Section 8 discusses the results of the proposed approach for three simulation scenarios.
2 Related Work
We explore the literature on optimizing sensor trajectories using the FIM of the likelihood function of the range measurements, where the likelihood function parameters correspond to the target positions. Various forms of the FIM establish a fundamental estimation lower bound limit for the mean squared error covariance matrix [26]. Therefore, many authors construct an objective as the D/A-optimality [27] of the PFIM, parameterized by sensor and estimated target positions [28]. We divide these works into two categories: those that consider static-sensors versus those that consider dynamic-sensors.
2.1 Static Sensors
Static sensors lack dynamic models or control inputs. Optimizing their placement involves determining where they would be positioned if placed instantaneously. Many studies on optimal radar placement use range measurements assumed to follow Gaussian distributions to construct the FIM, where the covariance may or may not vary with sensor-target distance [29, 13, 14, 16, 30, 31].
Several studies (e.g., [29, 13]) suggest that when the range measurement covariance is distance-independent, optimal sensor placement depends only on the angle between the sensor and the target. However, this approach overlooks the impact of signal path attenuation, which ideally results in higher information gain when the sensor is closer to the target.
To address this, [14, 32] adopt a heuristic assumption concerning covariance, combining a constant diagonal matrix with another diagonal matrix which is a function of the range between sensors and targets. However, signal processing fundamentals provide more realistic range measurement distributions. We take a similar approach to [16, 31], where the covariance, which scales with the return radar power, is determined by optimizing the Taylor series expansion of a match filter over time delay for a Narrowband return signal subject to attenuation and passing through an Additive White Gaussian Noise (AWGN) channel.
2.2 Dynamic Sensors
Practically, radars cannot simply “appear” at the optimal location; they require constrained dynamics and path planning. Moreover, dynamic sensors are essential to track dynamic targets, as the optimal radar configuration at one time step may not remain optimal at the next time step.
Historically, two main approaches have been used for optimal radar placement in target localization of dynamic targets: 1) moving the radars to track dynamic targets while maintaining an optimal radar-target localization geometry [33, 4], and 2) “spraying” radars over a large area to ensure high spatial coverage and multiplicity, then selecting which radars to employ for sensing [15]. This paper primarily focuses on the first approach, which is more appropriate given a limited number of radars.
Since the objective is a function of a finite horizon of time steps, many studies utilize an MPC [17] objective for optimizing the control inputs of mobile sensors [18, 19, 20, 21, 22]. Consequently, rather than directly optimizing the radar’s position at a specific time step, as in the static-sensor formulation, the optimal radar trajectory is indirectly determined by optimizing the control inputs to the radar kinematic model using the method of direct shooting [34].
The finite horizon MPC objective, formulated in terms the PFIM, may be efficiently computed through recursive evaluation using discrete-time nonlinear filtering [12]. Several studies on cooperative range-based underwater target localization with multiple autonomous surface vehicles adopt this approach, while also considering constraints such as distances between targets and sensors [9, 22, 21]. However, these methods often assume a constant diagonal covariance matrix for range measurements, irrespective of range, typically involve slow-moving sensors and targets, assume the target remains at a constant depth, and may assume a zero covariance for the target dynamic model. Moreover, solving the resulting nonlinear optimization problems remains challenging, with several recent approaches leveraging CasADi [35].
He et al [36] employs a more complex measurement model which integrates Multi-Target Tracking (MTT) characteristics such as data association, track insertion, and track deletion using the Joint Probabilistic Data Association Filter (JPDAF). However, this work assumes a constant diagonal covariance matrix for range and bearing measurements. Furthermore, the trajectory optimization only optimizes the sensor’s heading angle one time step ahead due to the highly nonlinear objective function, potentially leading to myopic and suboptimal trajectories.
Similar to information-based MPC objectives like the D/A optimality of the FIMs, [37] minimizes the mutual information gain between the target state distributions before and after the update step of a linear Kalman Filter, while parameterizing the measurement noise covariance based on the robot’s state. However, the measurement covariance only scales linearly with the sensor-target distance and is clipped at an extremely small variance of 0.01 meters, restricting experiments to small square arenas of 40 to 60 meters.
Hence, the proposed approach aims to enhance target localization through continuously optimizing radar placement by integrating an improved range measurement model into the objective function. Additionally, we leverage MPPI control to increase flexibility in solving the MPC objective over non-smooth and highly nonlinear objectives, enabling real-time control capabilities.
3 Preliminaries
We consider $N$ radars and $M$ targets. The state for radar $n$ at time step $k$ is denoted as $\bm{\chi}^{R_{n}}_{k}=[x,y,z,\theta,v,\omega]$, while for target $m$, it is $\bm{\chi}^{T_{m}}_{k}=[x,y,z,\dot{x},\dot{y},\dot{z}]$, where $\theta$ is the radar’s heading angle with respect to the $x$-axis of the radar coordinate frame, $v$ is the heading velocity, and $\omega$ is the angular velocity. The positions for radar $n$ and target $m$ are defined as $\bm{\chi}^{R_{n}}_{xyz}=[x,y,z]$ and $\bm{\chi}^{T_{m}}_{xyz}=[x,y,z]$, respectively. The time step $k$ has a duration of $\Delta t$. We denote the Hadamard product (elementwise multiplication) as $\odot$, and $\otimes$ is the kronecker product. Additionally, we use the notation $\Delta_{x}^{y}=\nabla_{x}\nabla_{y}^{T}$ for the Hessian operator.
To highlight the focus of our paper, we make the following simplifying assumptions to our analysis:
  1. 1.
    Perfect data association and target detection. We assume that the process of selecting which measurement(s) to incorporate into the target state estimator, is perfect, and the radars constantly track the targets.
  2. 2.
    Doppler shift can be neglected. For slow relative velocity $\nu$ between the radar and target, when the Doppler shift is much smaller than the effective bandwidth of the return signal, $\frac{f_{d}}{\beta}\ll 1$, the Doppler shift $f_{d}=2\times\frac{f_{c}}{c}\left(\frac{(\bm{\chi}^{T}_{xyz}-\bm{\chi}^{R}_{xyz})^{T}}{|\bm{\chi}^{T}_{xyz}-\bm{\chi}^{R}_{xyz}|}\right)\nu$ [38] may be disregarded in the radar return signal. $f_{c}$ is the carrier frequency and $c$ is the speed of light.
  3. 3.
    Radars have zero altitude. The targets move in $R^{3}$, but the radars move on a plane at ${z=0}$. This is a common assumption that radars are attached to ground vehicles.
  4. 4.
    We assume targets follow a constant velocity model with acceleration process noise and therefore do not account for drastic target maneuvers in our model.
With these assumptions, the following sections detail each computation block needed to continuously optimize radar placement (flow diagram depicted in Figure 2). We first discuss the more realistic range measurement model based on [16] compared to previous works such as [9, 22, 13].
Figure 2: Block Diagram of proposed approach.
4 Radar Signal Model
Following the approach in [16], we assume there are $N$ radars transmitting orthogonal narrowband signals to track $M$ targets. In this setup, the received signal at radar $n$ for target $m$ is given by:
$r^{nm}(t)$ $=\sum_{b=1}^{N}\zeta^{nm}e^{-j2\pi f_{c}\tau^{nm}}s^{b}(t-\tau^{nm})+\epsilon_{a}^{n}(t),$ (1)
where the time delay is:
$\tau^{nm}$ $=\frac{\lVert\bm{\chi}^{R_{n}}_{xyz}-\bm{\chi}^{T_{m}}_{xyz}\rVert+\lVert\bm{\chi}^{R_{b}}_{xyz}-\bm{\chi}^{T_{m}}_{xyz}\rVert}{c},$ (2)
and the set of orthonormal lowpass equivalent signals $s(t)$ such that there is no signal interference when multiple radars transmit simultaneously. The return signal noise $\epsilon_{a}(t)$ is circularly symmetric, zero-mean, complex Gaussian noise with autocorrelation function $\sigma_{a}^{2}\delta(\tau)$. $\zeta$ is the amplitude of the return signal, where $|\zeta|^{2}$ is the radar power return specified by the radar equation [39]:
$|\zeta|^{2}=P_{r}=\frac{P_{t}\Lambda_{t}\Lambda_{r}\lambda^{2}\Xi}{(4\pi)^{3}\lVert\bm{\chi}^{R}_{xyz}-\bm{\chi}^{T}_{xyz}\rVert^{4}L}$ (3)
The radar parameters specifying the received power $P_{r}$ $[W]$ at the radar include $P_{t}$ as the transmit power $[W]$, $\Lambda_{t}$ and $\Lambda_{r}$ as the transmit and receive gains, $L$ as the general loss factor, $\Xi$ as the radar cross section $[m^{2}]$, and $\lambda$ as the carrier frequency wavelength $[m]$.
It has been shown that the range measurement between radar $n$ and target $m$ approximately follows a Gaussian Probability Density Function (PDF) with variance parameterized by the radar-target geometry [13].
$z^{nm}\sim\mathcal{N}\left(2\lVert\bm{\chi}^{R_{n}}_{xyz}-\bm{\chi}^{T_{m}}_{xyz}\rVert\;,\;\frac{c^{2}\sigma_{a}^{2}}{8\pi^{2}f_{c}^{2}|\zeta^{nm}|^{2}}\right)$ (4)
We denote all the radar parameters, signal parameters, and noise parameters as a constant $\Gamma=\frac{\sigma_{a}^{2}\pi L}{2P_{t}\Lambda_{t}\Lambda_{r}\Xi}$. For brevity, we define shorthand notation as $\mu_{z^{nm}}=2\lVert\bm{\chi}^{R_{n}}_{xyz}-\bm{\chi}^{T_{m}}_{xyz}\rVert$ and $\sigma_{z^{nm}}=\Gamma\lVert{\bm{\chi}^{R_{n}}_{xyz}-\bm{\chi}^{T_{m}}_{xyz}}\rVert^{4}$.
Existing research has explored optimal radar placement when the covariance matrix of the range measurement depends on the radar-target geometry [14, 31, 40], primarily focusing on stationary targets and radars. Given that the range measurement model relies on the positions of both the radars and targets, the next section describes the target and radar kinematic models. This will facilitate the definition of the targets’ conditional transition probability in the targets’ state space model.
5 Process Model
We describe the kinematic model of a target to finalize the probabilistic state space model for the concatenated state of all the targets, encompassing both a conditional transition probability and the conditional measurement probability described in 4.
5.1 Target Kinematics: Constant Velocity Model
The target kinematic model for a single target follows constant velocity motion with acceleration noise [41, 42], which specifies the single-target transition model:
$\bm{\chi}^{T}_{k+1}$ $=A_{\text{single}}\bm{\chi}_{k}^{T}+\epsilon_{w}$ (5)
where
$A_{\text{single}}$ $=\begin{bmatrix}\bm{I}_{3\times 3}&\Delta t~{}\bm{I}_{3\times 3}\\ \mathbf{0}_{3\times 3}&\bm{I}_{3\times 3}\end{bmatrix}$ (6)
and $\epsilon_{w}$ is the acceleration noise which follows a zero mean Gaussian PDF with covariance
$\bm{W}_{\text{single}}$ $=\bm{W}_{\Delta t}\mathbf{\Sigma}_{w}\bm{W}_{\Delta t}^{T}$ (7)
where
$\bm{W}_{\Delta t}=\begin{bmatrix}\frac{1}{2}\Delta t^{2}~{}\bm{I}_{3\times 3}\\ ~{}\Delta t~{}~{}\bm{I}_{3\times 3}\end{bmatrix}~{},~{}\Sigma_{w}$ $=\bm{I}_{3x3}~{}\sigma_{W}^{2}$ (8)
and $\sigma_{W}$ determines the degree of deviation of the dynamics from constant velocity via the strength of the acceleration variance. Thus, the targets’ transition model (for all targets concatenated as a column vector) is:
$\bm{\chi}^{T_{1:M}}_{k+1}$ $=A\bm{\chi}^{T_{1:M}}_{k}+\epsilon_{W}$ (9)
$A$ $=\bm{I}_{M}\otimes A_{\text{single}}\quad W=\bm{I}_{M}\otimes W_{\text{single}}$ (10)
where $I_{M}\in\Re^{M\times M}$ is the identity matrix. Similarly, the measurements of all radar-target pairs may be concatenated to measurement measurement model of the targets’ state-space equations (based on Equation 4):
$\mathbf{z}^{T_{1:M}}$ $=\begin{bmatrix}\mu_{z^{11}}&\ldots&\mu_{z^{1M}}\\ \mu_{z^{21}}&\ldots&\mu_{z^{2M}}\\ \vdots&\ddots&\vdots\\ \mu_{z^{N1}}&\ldots&\mu_{z^{NM}}\\ \end{bmatrix}+\begin{bmatrix}\sigma_{z^{11}}&\ldots&\sigma_{z^{1M}}\\ \sigma_{z^{21}}&\ldots&\sigma_{z^{2M}}\\ \vdots&\ddots&\vdots\\ \sigma_{z^{N1}}&\ldots&\sigma_{z^{NM}}\\ \end{bmatrix}\odot\epsilon_{\mathbf{z}}$ (11)
$=\mathbf{H(\bm{\chi}^{T_{1:M}}_{xyz}})+\Omega\odot\epsilon_{\mathbf{z}}$ (12)
where $\sigma_{z^{nm}}$ is the standard deviation of the range measurement for radar $n$ and target $m$, as defined by the variance in Equation 4, while $\epsilon_{\mathbf{z}}$ is standard normal Gaussian PDF. Therefore the final state-space equations for the targets is shown in Equations 14 and 13:
$\bm{\chi}^{T_{1:M}}_{k+1}$ $=A\bm{\chi}^{T_{1:M}}_{k}+\epsilon_{W}$ (13)
$\mathbf{z}^{T_{1:M}}_{k+1}$ $=\mathbf{H}\left((\bm{\chi}^{T_{1:M}}_{xyz})_{k+1}\right)+\Omega\odot\epsilon_{\mathbf{z}}$ (14)
Given our focus on dynamic targets and dynamic radars, it’s essential to outline a basic kinematic model where the radar position is determined by control inputs and adheres to specific control and dynamic constraints. Thus, we next define the second-order unicycle kinematic model for the radars.
5.2 Radar Kinematics: Second-Order Unicycle Model
Each radar follows the second-order unicycle model. We use the Euler discretization of the unicycle continuous dynamic model with heading acceleration and angular acceleration control inputs $u=[u_{a},u_{\dot{\omega}}]^{T}$ to define the discrete time kinematic model:
$\bm{\chi}^{R}_{k+1}$ $=\bm{\chi}^{R}_{k}+G_{k}(\bm{\chi}^{R}_{k},u_{k})$ (15)
with control, velocity, and angular velocity limits,
[…]
With the radars’ positions defined relative to control inputs, the following section outlines the derivation of the FIMs with respect to the targets’ positions. This is based on the radar-target measurement model outlined in Equations 4 and 11, and the radars’ positions parameterized by Equations 15 and 16.
6 Fisher Information Matrix
The Standard Fisher Information Matrix (SFIM) measures the amount of information that an observable random variable $z$ (range measurements) carries about an unknown parameter $\theta$ (targets’ states) [12, 43]:
$\bm{J}_{D}(\theta)=E\left[\left(\nabla_{\theta}\log f\left(z|\theta\right)\right)\left(\nabla_{\theta}\log f\left(z|\theta\right)\right)^{T}\right]$ (17)
Under the assumption of independence, the range measurements likelihood in Equation 11 is represented as:
$f\left(\mathbf{z}^{T_{1:M}}|\bm{\chi}^{T_{1:M}}_{xyz};\bm{\chi}^{R_{1:N}}_{xyz}\right)$ $=\prod_{n=1}^{N}\prod_{m=1}^{M}f(z^{nm}|\bm{\chi}^{T_{m}}_{xyz};\bm{\chi}^{R_{n}}_{xyz})$ (18)
Given that the measurement is a Gaussian PDF, with the mean and covariance parameterized by the targets’ positions, and that information of measurements is additive by the independence assumption, we use the derivation of general Gaussian FIM [44] to calculate the SFIM for a multiple targets’ positions with multiple radars:
$\bm{J}(\bm{\chi}^{T_{m}}_{xyz};\bm{\chi}^{R_{n}}_{xyz})$ $=(\bm{\chi}^{T_{m}}_{xyz}-\bm{\chi}^{R_{n}}_{xyz})(\bm{\chi}^{T_{m}}_{xyz}-\bm{\chi}^{R_{n}}_{xyz})^{T}\times$
$\left(\frac{4}{\Gamma\lVert\bm{\chi}^{R}_{xyz}-\bm{\chi}^{T}_{xyz}\rVert^{6}}+\frac{8}{\lVert\bm{\chi}^{R}_{xyz}-\bm{\chi}^{T}_{xyz}\rVert^{4}}\right)$ (19)
$\bm{J}(\bm{\chi}^{T_{m}}_{xyz};\bm{\chi}^{R_{1:N}}_{xyz})$ $=\sum_{n=1}^{N}\bm{J}(\bm{\chi}^{T_{m}}_{xyz};\bm{\chi}^{R_{n}}_{xyz})$ (20)
$\bm{J}(\bm{\chi}^{T_{1:m}}_{xyz};\bm{\chi}^{R_{1:n}}_{xyz})=$
$\begin{bmatrix}\bm{J}(\bm{\chi}^{T_{1}}_{xyz};\bm{\chi}^{R_{1:N}}_{xyz})&\!\!0&\!\!\ldots&\!\!0\\ 0&\!\!\bm{J}(\bm{\chi}^{T_{2}}_{xyz};\bm{\chi}^{R_{1:N}}_{xyz})&\!\!\ldots&\!\!0\\ \vdots&\!\!\vdots&\!\!\ddots&\!\!\vdots\\ 0&\!\!0&\!\!\ldots&\!\!\bm{J}(\bm{\chi}^{T_{M}}_{xyz};\bm{\chi}^{R_{1:N}}_{xyz})\end{bmatrix}$ (21)
Equation Equation 19 is the SFIM of a single target position parameterized by a single radar and Equation 20 is the SFIM of a single target positions parameterized by multiple radars.
When there is prior knowledge about the dynamics of the target state (Equation 5), then this information may be incorporated into the SFIM to form the PFIM:
$\bm{J}=\bm{J}_{D}+\bm{J}_{P}$ (22)
$\bm{J}_{P}=E[\Delta_{\bm{\chi}^{T}}^{\bm{\chi}^{T}}\log f(\bm{\chi}^{T})]$ (23)
Here, $\bm{J}_{D}$ represents the information about the unknown target state obtained from the data, while $\bm{J}_{P}$ reflects the information about the unknown target state based on prior assumptions. Now, considering that the targets’ state includes velocity, we update Equation 21 to be $\bm{J}_{D}$ for the PFIM, with the block matrices as:
$\bm{J}(\bm{\chi}^{T_{n}};\bm{\chi}^{R_{1:N}}_{xyz})=\begin{bmatrix}\bm{J}(\bm{\chi}^{T_{n}}_{xyz};\bm{\chi}^{R_{1:N}}_{xyz})&0\\ 0&0\end{bmatrix}$ (24)
For discrete time filtering systems, such as Equations 14 and 13, the Equation 23 is recursively computed at time $k+1$ from time step $k$ using the famous recursion updates in [12]. The recursion equations are written as
$\bm{J}_{k+1}(\bm{\chi}^{T_{1:M}}_{k+1})=\bm{D}_{k}^{22}-\bm{D}_{k}^{21}(\bm{J}_{k}+\bm{D}_{k}^{11})^{-1}\bm{D}_{k}^{12}$ (25)
where
$\bm{D}_{k}^{11}=E[-\Delta_{\bm{\chi}^{T_{1:M}}_{k}}^{\bm{\chi}^{T_{1:M}}_{k}}\log p(\bm{\chi}^{T_{1:M}}_{k+1}|\bm{\chi}^{T_{1:M}}_{k})]$ (26)
$\bm{D}_{k}^{12}=E[-\Delta_{\bm{\chi}^{T_{1:M}}_{k}}^{\bm{\chi}^{T_{1:M}}_{k+1}}\log p(\bm{\chi}_{k+1}|\bm{\chi}_{k})]=[\bm{D}_{k}^{21}]^{T}$ (27)
$\bm{D}_{k}^{22}=E\left[-\Delta_{\bm{\chi}^{T_{1:M}}_{k+1}}^{\bm{\chi}^{T_{1:M}}_{k+1}}\log p(\bm{\chi}^{T_{1:M}}_{k+1}|\bm{\chi}^{T_{1:M}}_{k})\right]+$
$\qquad E\left[-\Delta_{\bm{\chi}^{T_{1:M}}_{k+1}}^{\bm{\chi}^{T_{1:M}}_{k+1}}\log p(z_{k+1}|\bm{\chi}^{T_{1:M}}_{k+1})\right]$ (28)
Each $D$ term in equation Equation 25 is (PDF regularity assumptions) as
$\bm{D}_{k}^{11}=\bm{A}^{T}\bm{W}^{-1}\bm{A}$ (29)
$\bm{D}_{k}^{12}=-\bm{A}^{T}\bm{W}^{-1}$ (30)
$\bm{D}_{k}^{22}=\bm{W}^{-1}+E_{f(\bm{\chi}^{T_{1:M}}_{1:k+1})}\left[\bm{J}_{\bm{D}_{k}}(\bm{\chi}^{T_{1:M}}_{k+1}))\right]$ (31)
where due to the Markov assumption on the transition model in the targets’ state space equation (Equation 9), the joint PDF $f(\bm{\chi}^{T_{1:M}}_{1:k})$ is simplified as
$f(\bm{\chi}^{T_{1:M}}_{1:k})=f(\bm{\chi}^{T_{1:M}}_{0})\prod_{t=1}^{k+1}f(\bm{\chi}^{T_{1:M}}_{t}|\bm{\chi}^{T_{1:M}}_{t-1})$ (32)
After substituting equations Equations 29, 30 and 31 into Equation 25, and leveraging the Woodbury matrix identity [26], we arrive at [45]:
$\bm{J}_{k+1}(\bm{\chi}^{T_{1:M}}_{k+1})$ $=(\bm{W}+\bm{A}^{T}\bm{J}_{k}^{-1}\bm{A})^{-1}+$
$\quad\quad E_{f(\bm{\chi}^{T_{1:M}}_{1:k+1})}\left[\bm{J}_{\bm{D}_{k}}(\bm{\chi}^{T_{1:M}}_{k+1}))\right]$ (33)
Equation 33 is preferred over Equation 25 for numerical stability. This preference arises because it avoids the inversion of the $W$ matrix, which can be close to singular in cases where the conditional transition PDF is described by constant velocity models.
To compute the expectation over the target state for Equation 31, as well as objective terms involving Equation 21 or any other objective term in the MPC objective that involves the targets’ states, we employ the Cubature Kalman Filter (CKF) to describe Equation 32.
CKFs efficiently leverage third-order spherical cubature integration into both the prediction and correction steps of an additive Gaussian noise Kalman filter [46]. To compute the expectations described in the previous paragraph, we propagate the sigma points of the CKF multiple time steps into the future, employing a number of sigma points equal to twice the dimension of the state variable. Compared to a Particle Filter (PF) [46], the CKF has lower computational complexity, as the number of evaluation points scales linearly with the state size, while the number of particles in a PF may substantially increase.
With our current formulations in Sections 4, 5 and 6, we formulate a discrete-time finite-horizon MPC objective using dynamic, control, and mission-aware constraints.
7 Model Predictive Controller Objective
MPC is an optimal control technique that determines the control actuation minimizing a cost function for a constrained dynamical system over a finite, receding horizon [47, 48]. At each time step of the receding horizon objective, we enforce minimum distances between radars and targets to prevent potential destruction by adversaries using soft penalties. Similarly, we ensure safe separation distances between mobile radars to prevent collisions. Additionally, we incorporate the radar’s acceleration, angular acceleration, angular speed, and heading velocity constraints as described in Equation 16 into the dynamic equation using clipping functions. Finally, we directly integrate the radar’s dynamic equation (Equation 16) into the objective using direct shooting [34].
As a result, converting hard constraints to soft/implicit constraints results in an unconstrained, discontinuous, and non-differentiable optimization formulation. Lastly, we incorporate the uncertainty of the targets’ state by taking expectation of the MPC objective function with respect to the CKF estimator of the targets’ state:
$S(\bm{u}_{k:k+K+1};\bm{\chi}^{T_{1:M}}_{k:k+K+1})\!=\!E_{\tilde{f}}\bigr{[}S_{traj}(\bm{u}_{k:k+K+1};\bm{\chi}^{T_{1:M}}_{k:k+K+1})$
$\quad\quad\quad\quad\quad\quad\quad+\alpha_{1}S_{R2R}(\bm{u}_{k:k+K})$
$\quad\quad\quad\quad\quad\quad\quad+\alpha_{2}S_{R2T}(\bm{u}_{k:k+K};\bm{\chi}^{T_{1:M}}_{k:k+K})\bigr{]}$ (34)
where
$S_{traj}\left(\bm{\chi}^{R_{1:N}}_{k:k+K+1};\bm{\chi}^{T_{1:M}}_{k:k+K+1}\right)=$
$\quad\sum_{t=k}^{k+K+1}\gamma^{t-k}\log\det\left(J(\bm{\chi}_{t}^{T_{1:M}};\bm{\chi}_{t}^{R_{1:M}}\right)$
$S_{R2T}(\bm{u}_{k:k+K+1}\!)\!=$
$\quad\;\;\!\!\sum_{t=k}^{k+K+1}\!\!\gamma^{t-k}\!\!\sum_{m=1}^{M}\!\sum_{n=1}^{N}\!\mathbf{1}\!\left(\lVert\bm{\chi}_{t}^{T_{n}}\!-\!\bm{\chi}_{t}^{R_{m}}(\bm{u}_{t})\rVert\!\leq\!R_{R2T}\!\right)$
$S_{R2R}(\bm{u}_{k:k+K+1})=$
$\quad\;\sum_{t=k}^{k+K+1}\gamma^{t-k}\!\!\!\!\!\!\sum_{\{i<j:i,j\in M\}}\!\!\!\!\!\!\mathbf{1}\left(\lVert\bm{\chi}_{t}^{R_{i}}(\bm{u}_{t})-\bm{\chi}_{t}^{R_{j}}(\bm{u}_{t})\rVert\leq R_{R2R}\right)$ (35)
where $\alpha_{1},\alpha_{2}$ are hyper parameters that determine how much to penalize the radars for violating the soft constraints $S_{R2R},S_{R2T}$ respectively, $\gamma$ is the discount factor, and $K$ is the horizon length. The CKF estimator PDF on the targets’ state over the horizon is
$\tilde{f}(\bm{\chi}_{k:k+K+1}^{T_{1:M}})=f_{ckf}(\bm{\chi}^{T_{1:M}}_{k}|\mathbf{z}_{1:k-1})\!\prod_{t=k+1}^{k+K+1}\!f(\bm{\chi}^{T_{1:M}}_{t}|\bm{\chi}^{T_{1:M}}_{t-1})$ (36)
This MPC problem requires an efficient solver that can handle highly nonlinear, discontinuous, and non-differentiable functions. Therefore, we employ a sampling-based strategy, MPPI [24], as discussed in the following section.
7.1 Model Predictive Path Integral
MPPI is a sampling-based MPC algorithm designed for optimizing nonlinear stochastic systems subject to complex cost functions in real time [49]. The assumption of MPPI is the control input is random and distributed as Gaussian random vector $\bm{v}\sim N(\bm{u},\bm{\Sigma})$. Due to the stochasticity of the control input, rather than solving for a deterministic optimal control input we solve for an optimal control PDF $Q^{*}$ [50]:
$Q^{*}=\frac{1}{\eta}\exp\left(-\frac{1}{b}S(\bm{v}_{k:k+K+1})\right)P(\bm{v}_{k:k+K+1})$ (37)
where $P(\cdot)$ is the PDF of an control input sequence in an uncontrolled system ($\bm{u}_{k:k+K+1}=0)$, $b$ is the temperature of the distribution, and $\eta$ is the evidence. However, the optimal control PDF likely does not possess an analytical form. Therefore, by minimizing the Kullback-Leibler Divergence (KL-Divergence) between $Q^{*}$ and a variational Gaussian PDF $Q$ parameterized by $u,\Sigma$ [24] with respect to $u$ gives the optimal control at $k$ as [24]
[…]
Since we cannot express $Q^{*}$ analytically in Equation 38, we leverage AIS to approximate the expectation as follows [25, 24]:
  1. 1.
    sample control trajectories $\bm{v}_{k:k+K+1}^{(i)}$ from distribution $Q$ with mean $\bm{u}_{k:k+K+1}$ and covariance $\Sigma$
  2. 2.
    evaluate the objective function for each control trajectory sample $S(\bm{v}_{k:k+K+1}^{(i)})$
  3. 3.
    compute the importance sampling corrective weights for each sample $\tilde{w}^{(i)}\propto S(\bm{v}^{(i)})$
  4. 4.
    adapt the mean and covariance of the PDF $Q$
  5. 5.
    repeat steps 1-4
Due to the high dimensionality of the targets’ control input over a horizon and the limited number of samples in MPPI, the sample covariance matrix for control inputs can overfit, leading to inaccurate variances [51]. This often results in poor control inputs from the proposal distribution of MPPI. To address this, we adopt the Ledoit-Wolf (LW) for AIS, which approximates Oracle Approximating Shrinkage (OAS) shrinkage and helps mitigate overfitting, following the method described in [25]  [^1] The authors in [25] do not explicitly discuss Oracle Approximating Shrinkage (OAS) or Ledoit-Wolfe, but their publicly available code demonstrates methods for estimating the sample covariance, including Ledoit & Wolf, Schaffer & Strimmer, Rao-Blackwell, and OAS. A reference in the publicly available code is https://github.com/sisl/MPOPIS/wiki/MPOPIS-Details..
(a) Proposed approach: Each radar moves towards the targets.
(b) [22, 13, 9] range model: The upper right radar moves away from targets
(c) Radars move toward targets as closely as feasible.
(d) Radars encircle targets, but the bottom most radar remains distant.
(e) Radars maintain circular formation with targets circle formation.
(f) Radars encircle the targets, but bottom most radar moves away from targets.
(g) Radars maintain a circular formation within the circle of targets.
(h) Radars maintain encirclement formation around targets cricle
Figure 3: The left column shows snapshots of target localization using a range measurement model with covariance dependent on radar-target distance [16], while the right column depicts tracking with a range measurement model employing constant diagonal covariance [13]. In the first three rows, each subfigure’s left panel shows the 3D projection onto 2D of continuously optimizing radar placement at a specific time. The colored lines emanating from the radars depict the weighted MPPI-planned trajectories, with blue indicating better paths and red indicating worse ones. Dashed red and green circles mark the radar’s (10 [m]) and the target’s (125 [m]) collision avoidance boundaries, respectively, while blue dashed circles represent the targets’ collision avoidance boundary based on the CKF position estimate. The radar’s trajectory is the last 25 time steps. The right panel displays the log determinant of the FIM (for the ddr measurement model) concerning the 2D target position for the specific time step, meshed over the free space area, and evaluated at each mesh point.
(a) 3 Radar - 4 Target: RMSE with 90% HDI
(b) 3 Radar - 4 Target: RMSE without stationary radar or HDI
(c) 3 Radar - 4 Target: RMSE ECDF
(d) 6 Radar - 3 Target: RMSE with 90% HDI
(e) 6 Radar - 3 Target: RMSE without stationary radar or HDI
(f) 6 Radar - 3 Target: RMSE ECDF
(g) 4 Radar - 4 Target: RMSE with 90% HDI
(h) 4 Radar - 4 Target: RMSE without stationary radar or HDI
(i) 4 Radar - 4 Target: RMSE ECDF
Figure 4: The difference in target localization errors between constant covariance and radar-based covariance for 3 radars aand 4 targets, 6 radars aand 3 targets, and 4 radars and 4 targets for the SFIM and PFIM based objective via RMSE plot with 90% HDI and RMSE ECDF plots. The left column shows the mean RMSE over the targets’ true state versus the CKF predicted state over 500 MC trials. The lightly shaded region is the 90% highest density interval (HDI) of the RMSE over 500 MC trials. The middle column shows a zoom in of the left column subfigurese without stationary radar and 90% HDI RMSEs. The right column is the empirical ECDF of the CKF predicted state RMSE over time steps and MC trials. Higher ECDF is better at a given RMSE.
Data: $\bm{\chi}^{R_{1:N}}_{0}$
Result: $\hat{\bm{\chi}}^{T_{1:M}}_{1:T},\bm{\chi}^{R_{1:N}}_{1:T}$
for $k=1:T$ do
       $\bm{\chi}^{T_{1:M}}_{k+1}$ = TransitionFn$\left(\bm{\chi}^{T_{1:M}}_{k}\right)$
       $\hat{\bm{\chi}}^{T_{1:M}}_{k+1}$ = CKF.predict$\left(\hat{\bm{\chi}}^{T_{1:M}}_{k}\right)$
       if $k\%T_{control}=0$ then
             $\hat{\bm{\chi}}^{T_{1:M}}_{k+1:k+K+1}\leftarrow$ CKF.propogate$\left(\hat{\bm{\chi}}^{T_{1:M}}_{k+1}\right)$;
[…]
       end if
      $\bm{\chi}^{R_{1:N}}_{k+1}\leftarrow$ Actuate$\left(\bm{\chi}^{R_{1:N}}_{k},\mathbf{u}_{k}\right)$ ;
       $\mathbf{z}^{R_{1:N}}_{k+1}$ = MeasureFn$\left(\bm{\chi}^{T_{1:M}}_{k+1},\bm{\chi}^{R_{1:N}}_{k+1}\right)$ ;
       $\hat{\bm{\chi}}^{T_{1:M}}$ = CKF.update$\left(\bm{\chi}_{k+1}^{R_{1:N}},\mathbf{z}^{R_{1:N}}_{k+1}\right)$ ;
end for
Algorithm 1 Continuous Optimal Radar Placement
The integration of the algorithms outlined in Sections 4, 5, 6 and 7 forms the enhanced continuous optimal radar placement pipeline, described in Algorithm 1. This pipeline includes two key enhancements: an improved range measurement model, which significantly impacts the FIM, and a flexible and computationally efficient MPC objective optimizer using MPPI.
8 EXPERIMENTS
We examine three scenarios: 1. an “underdetermined” system, with fewer radars than targets, 2. an “overdetermined” system, with more radars than targets, and 3. an “full” system, with equal amount of radar and targets. We highlight not only the poorer performance of a misspecified measurement model (DDR versus constant covariance range (CCR) measurement model) but also the pathologies in radar movement caused by the controller using a FIM-based objective specified from the CCR measurement model. The proposed approach outperforms methods employing stationary radars and simplified range measurement models in target localization, achieving a reduction in mean RMSE and the upper tail of the 90% HDI by 38-74% and 33-79% respectively, across 500 Monte Carlo (MC) trials at all time steps.
8.1 Simulation
The hyperparameters for each experiment for the MPPI control, the AIS, the radar configurations , and the MPC objective is found in Tables 3, 2 and 1.
Radar Parameter Value
Carrier Frequency $1\times 10^{8}$
Radar Loss 1
Transmitter Gain 200
Receiver Gain 200
Radius (for SNR) 500 [m]
Radar Cross Section 1 [m2]
Power Transmit 1000 [W]
SNR -20 [dB]
Table 1: The Radar Equation configuration settings for each experiment
MPPI Parameters Value
$u_{a}$ std 25 [m/s2]
$u_{\dot{w}}$ std $45^{\circ}$
Horizon 15
Number of Sample Trajectories 200
Number of MPPI Sub-iterations 5
Temperature 0.1
Elite Threshold 0.9
AIS Method Cross Entropy
Covariance Shrinkage Method Ledoit-Wolf Shrinkage
Radar Cross Section 1
Table 2: The MPPI settings for each experiment
MPC Parameter Value
$\gamma$ $0.95$
$R2T$ 125 [m]
$R2R$ 10 [m]
$\alpha_{1}$ 500
$\alpha_{2}$ 1000
Table 3: The MPC configuration settings for each experiment
We assume that the radars move on the ground plane (z=0), while the drones move in free space by the constant velocity kinematic model (Equation 5) with acceleration noise constant $\sigma_{W}=\sqrt{10}$. For each experiment , we conduct 500 MC trials, where every MC trial is 600 time steps (1 minute in real time). The initial heading velocity and acceleration of the radars are both set to 0, and their initial positions in the XY-plane are uniformly sampled within a square with edge sizes 800 [m], centered at the origin (0 [m], 0 [m]). Although the targets may start in the same position for each MC trial, the random acceleration noise in the transition model (Equation 9) leads to significant variations in the trajectories of the targets across trials. The observed range measurements at each time step are simulated using the more realistic range measurement model presented in Equation 4.
We compare our approach quantitatively and qualitatively with stationary radars. Additionally, using the MPPI controller for mobile radars, we compare the improved measurement model [16] to the CCR measurement model used in [29, 13, 9, 22]. This comparison not only empirically demonstrates the suboptimality of the constant covariance model as range measurements become noisier with increasing distance between the radar and the target (due to model mispecification), but also highlights the pitfalls of a controller using an objective based on a CCR measurement model. In the CCR measurement model, we use an “optimistic” diagonal covariance matrix with elements $R2T^{4}*\Gamma$ meters, where $R2T$ is the radar-target distance constraint.
For experimental purposes, we maintain constant noise power for the AWGN in Equation 1. We establish the desired SNR level at a specified distance using the formula:
$\sigma_{a}^{2}=\frac{M\tilde{P_{r}}}{10^{SNR/10}}$ (39)
Here, $\tilde{P}_{r}$ represents the radar received power at the specified distance 500 [m]. For all experiments we set SNR = -20 [dB].
We note that the MPPI controller was written in JAX, which on a NVIDIA GeForce RTX 2080 Ti runs at 26-27 Hz (realtime), despite not fully optimizing our code.
8.2 Improved Target Localization
In over 500 MC trials, we empirically show that our approach exhibits significantly lower RMSE for all time steps (see Figure 4 (a,b,d,e,g,h)) compared to both the stationary radar approach and the mobile radar approach with constant range measurement. Additionally, the 90% HDI interval in localization performance for the CCR measurement model is significantly larger compared to the DDR measurement model presented in [16] (Figure 4 (c,f,i)). The PFIM-based objective for both range measurement models has slightly higher targets’ state estimate RMSE compared to the SFIM-based objective.
However, Bayesian methods are notorious for overfitting, attributed to ensembling with model mispecification and the assumption of a single true model generating the data [52, 53]. We further examine the performance of the proposed approach inn the three simulation scenarios in the next subsections.
8.2.1 3 Mobile Radars and 4 Targets
We initialize the targets such that all target directional velocities point towards the upper right (northeast) in the world coordinate frame, and begin around the same position (see Figure 3):
$\chi^{T_{1:M}}_{0}=\begin{bmatrix}0&0&55&20&10&0\\ 15.4&15.32&70&15&20&0\\ 10&10&55&17&19&0\\ 20&20&45&6&8&0\end{bmatrix}$
where the first 3 columns are the 3D xyz coordinates, and the last 3 columns are the directional velocities.
Therefore, with fewer radars than targets, we anticipate a circular formation of radars near the targets when continuously optimizing radar placement, aiming to maintain nearest feasible proximity to the targets.
To illustrate the shortcomings of a controller using a CCR measurement model-based objective for radar placement, we showcase four time steps of a MC simulation realization, as depicted in Figure 3. For our proposed approach, at the beginning of the simulation, radars move towards targets, forming a circular arrangement within the nearest feasible proximity while adhering to radar-target constraints. As targets disperse, radars maintain this formation, but increased range measurement noise prompts a tighter circle, resulting in most targets being slightly outside the radar formation (left subfigure in Figure 3). In contrast, approaches with distance-independent noise maintain optimal angular spacing regardless of radar-target distance, often moving very far from the targets (Figure 3 (b,d,f,h)) and causing high localization errors (Figure 4 (a-i)). Counterintuitively, to achieve the optimal angle, mobile radars using the CCR measurement model move away from the targets, resulting in larger localization errors (Figure 3(b,f)). The target localization error over 1000 time steps for the extended realized MC simulation in Figure 3 is depicted in Figure 5.
Figure 5: RMSE for single simulation realization in Figure 3
8.2.2 6 Mobile Radars and 3 Targets
We initialize each target’s velocity direction to be different, with target 1 heading towards the upper right, target 2 towards the lower right, and target 3 towards the lower left:
$\chi^{T_{1:M}}_{0}=\begin{bmatrix}0&0&70&25&20&0\\ -100.4&-30.32&45&20&-10&0\\ 30&30&80&-10&-10&0\\ \end{bmatrix}$
where the first 3 columns are the 3D xyz coordinates, and the last 3 columns are the directional velocities.
The initial positions of the targets are depicted in Figure 1. Since the targets’ are spreading out over an open space, we hypothesize that the radars will send atleast 1 radar to follow each target, while keeping a reasonable distance to all other targets.
As the targets disperse in free space, the 6 radars coordinate to position 2 radars on either side of each target’s direction of motion. This strategy intuitively avoids maintaining a circular formation, which would place the radars too far from the targets, leading to noisy measurements and inaccurate target localization (Figure 1 (a-h)). Optimal radar placement remains relatively similar regardless of whether a distance-independent or distance-dependent range noise model is used, though the radars are positioned farther from the targets in the former case which leads to slightly higher localization error and 90% HDI interval (Figure 1 (c,d,g,h)). In particular, the upper-most radar in Figure 1(h) is quite far away from the upper right-most target. The extended MC simulation over 1000 time steps is shown in Figure 1 with corresponding target localization error depicted in Figure 6.
Figure 6: RMSE for single simulation realization in Figure 1
8.2.3 4 Mobile Radars and 4 Targets
We initialize two targets to move towards the upper right and two targets to move towards the lower left. The initial positions of the targets are:
$\chi^{T_{1:M}}_{0}=\begin{bmatrix}0&15&70&15&15&0\\ 40.4&15&70&15&15&0\\ -30&-15&45&-10&-10&0\\ 20&-15&45&-10&-10&0\\ \end{bmatrix}$
where the first 3 columns are the 3D xyz coordinates, and the last 3 columns are the directional velocities.
We hypothesize that the radars will pair off, with two radars following each group of two targets.
As the targets disperse in free space, the 4 radars coordinate to position 2 radars on either side of each pair of targets’ direction of motion. This strategy also avoids maintaining a circular formation, which would place the radars too far from the targets, leading to noisy measurements and inaccurate target localization (Figure 2 (a-h)). Optimal radar placement for the CCR measurement model based controller is quite different from the DDR measurement model based controller, where the former positions one of the radars extremely far away from all targets in the simulation (Figure 2 (b,d,f,i)) leading to high localization error. In particular, the upper-left most radar in Figure 2(h) is quite far all targets. The extended MC simulation over 1000 time steps is shown in Figure 2 with corresponding target localization error depicted in Figure 7.
Figure 7: RMSE for single simulation realization in Figure 2
(a) 3 radars and 4 targets. The uppermost target shows significant height error, while the leftmost target exhibits substantial position error in the xy plane.
(b) 6 radars and 3 targets. The bottom rightmost target exhibits significant height error, while the top rightmost target displays a position error in the xy plane.
(c) 4 radars and 4 targets. The right uppermost target exhibits significant height error.
Figure 8: Snapshots of each simulation scenario for stationary radars highlight the poor geometric layout of stationary radars for target localization.
8.2.4 Stationary Radars
Mobile radar methods, whether employing constant covariance or distance-dependent covariance in the range measurement model, outperform stationary radars (Figure 4 (a-i)). We present a qualitative example for each simulation scenario, illustrating the failure of stationary radar to track targets (Figure 8 (a-c)) and showing how radar positioning affects the information gain of measurements. As targets move farther away from the stationary radars, the noise in the range measurements increases drastically, resulting in higher localization errors.
9 CONCLUSION
We implemented two improvements to optimize radar placement for multi-target localization: a more realistic radar range measurement model and an efficient stochastic optimal controller (Model Predictive Path Integral (MPPI)) for handling non-smooth, non-convex objectives in real time. Our approach significantly improves the Cubature Kalman Filter (CKF) target state estimate Root Mean Squared Error (RMSE) compared to stationary radars. Additionally, we show the pitfalls of the constant covariance range (CCR) measurement model, which may increase target localization error by moving radars in the opposite direction of target movement to optimize angles rather than the radar-target distance. Overall, the proposed strategy outperforms stationary radars and simplified range measurement models in target localization, achieving a 38-74% reduction in mean RMSE and an 33-79% reduction in the upper tail of the 90% Highest Density Interval (HDI) over 500 Monte Carlo (MC) trials across all time steps. For future work, we plan to integrate occupancy grid maps into the collision cost function to account for static and dynamic obstacles, such as trees and buildings, and further improve the radar range measurement model with traditional Multi-Target Tracking (MTT) characteristics.